OpenStudy (anonymous):

derrive h(x)=from 2 to (1/x) arctan t dt

OpenStudy (anonymous):

derivative?

OpenStudy (primeralph):

Drive.

OpenStudy (anonymous):

find derrivative

OpenStudy (anonymous):

oh maybe \[\frac{d}{dx}\int_2^{\frac{1}{x}}\arctan(t)dt\]

zepdrix (zepdrix):

\[\large h(x)=\int\limits_2^{1/x}\arctan t \;dt\] Find \(\large h'(x)\) Is this your problem? :o

OpenStudy (primeralph):